# Arithmetic on Continued Fractions

### v1

@@ 3.1. Failure to Converge
(9) $[z_{0},z_{1},\cdots]={\frac{9xy+3x+3y+1}{4xy+2x+2y+1}}$
(10) $[z_{0},z_{1},\cdots]={\frac{49xy+21x+21y+9}{25xy+20x+10y+4}}$
(11) $[z_{0},z_{1},\cdots]={\frac{289xy+119x+119y+49}{144xy+60x+60y+25}}$
@@
The source of the problem is now clear: we should have $z=[z_{0},z_{1}]=[2,\infty]$, but no finite number of terms can ever tell us that $z_{0}=2$. It is always possible that one input is in fact less than $\sqrt{2}$, so it is always possible that $z<2$ and $z_{0}=1$.
Our approach to the problem (see [2] for an entirely different approach) is to separate the internal representation of a CF as a data structure from the mathematical notion of a sequence of integer terms. The internal representation will contain explicit terms (i.e. $z_{k}=n$) whenever such can be determined, but might also contain arbitrary bounds on the tail $[z_{k},z_{k+1}\cdots]$, such as the sequence of bounds we derived for $\sqrt{2}*\sqrt{2}$. This is a direct generalization, because the explicit term $z_{k}=n$ is equivalent to the bound
@@
Therefore, when a CF is constructed explicitly from a known sequence of terms $[z_{0},z_{1},\cdots]$, it will be represented internally as the sequence of half-open intervals
$[[z_{0},z_{0}+1),[z_{1},z_{1}+1),\cdots]\ .$
@@
At the opposite extreme, the internal representation for $\sqrt{2}^{2}$ will consist entirely of ever-tighter bounds on the first term, i.e. the sequence of intervals
$[[\frac{16}{9},\frac{9}{4}),[\frac{49}{25},\frac{100}{49}),[\frac{576}{289},\frac{289}{144}),\cdots]\ .$
@@
The general idea (elaborated more fully in section 3.3) is that we will generate output on every iteration of the arithmetic algorithm; when we cannot produce a term, we output upper and lower bounds on the current homomorphic matrix, leaving the matrix unchanged. Reading such output, there can be no ambiguity about the term to which a bound applies; the first bound applies to $z_{0}$, and as soon as we encounter a bound which uniquely determines $z_{k}$ (i.e. an interval with $\lfloor x\rfloor$ the same for all $x$ in the interval), we know that the next bound will apply to $z_{k+1}$. This implies that any finite prefix of the output sequence will determine a finite prefix $[z_{0}\cdots z_{k-1}]$ of the actual CF, followed by a bound $[\ell_{k},u_{k})$ on the tail $[z_{k},z_{k+1}\cdots]$. The exact value of the CF is between $[z_{0}\cdots z_{k-1},\ell_{k}]$ and $[z_{0}\cdots z_{k-1},u_{k}]$ (where we extend the usual CF notation by allowing a rational number, instead of an integer, as the last term).
@@
Note that storage of a long sequence of intervals is not an issue for implementation; successive bounds on the same term are always tighter, so there is no need to retain earlier bounds when a new one is obtained.
3.2. Extracting Terms
@@
We must note that our representation of CF arithmetic makes it impossible to simply ask a question like “what are the first five terms of $z=xy$?” An explicit list of terms can only be relative to a required degree of accuracy $\varepsilon$. Our arithmetic algorithms consume and return theoretically infinite sequences of intervals; extracting a finite sequence of terms is a separate “post-processing” step, carried out only when a numerical approximation is required. To obtain such an approximation, we find a prefix of the output interval sequence which is long enough to make the difference between $[z_{0}\cdots z_{k-1},\ell_{k}]$ and $[z_{0}\cdots z_{k-1},u_{k}]$ less than $\varepsilon$; then our approximate CF is $[z_{0}\cdots z_{k}]$, where $z_{k}$ is some integer in the interval $[\ell_{k},u_{k})$. We might as well always take $z_{k}=\lfloor u_{k}\rfloor$, which is the only integer in the interval if the interval is small.
In the case of $\sqrt{2}^{2}$, any accuracy threshold will yield the same result, $z=[2]$.
3.3. An Algorithm for CF Arithmetic
@@
We now present the full details of the algorithm suggested in the previous section. We augment the matrix $M$ with a pair of intervals $\{I_{x},I_{y}\}=\{[x_{\ell},x_{u}),[y_{\ell},y_{u})\}$, the current bounds on the remaining tails of the inputs $x,y$. These intervals start at (,)($\infty$,$\infty$)( - ∞ , ∞ ), since the first term could be any integer. When we ingest an interval of the form $[a,a+1)$ from $x$ or from $y$, we modify $M$ as before with $M\leftarrow\mbox{ingest\_x}(a,M)$ or $M\leftarrow\mbox{ingest\_y}(a,M)$, and we change that variable’s bound to $[1,\infty)$. When we ingest an ambiguous interval $[\ell,u)$ (i.e. a bound that does not determine the next term of the input), we replace the current bound on the corresponding variable with $[\ell,u)$, leaving $M$ unchanged.
We define $\mathcal{R}(M)$ to take the bounds into account, i.e.
(12) $\mathcal{R}\left(\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right),I_{x},I_{y}\right)=\left\{\left\lfloor{\frac{axy+bx+cy+d}{exy+fx+gy+h}}\right\rfloor|\ x\in I_{x},y\in I_{y}\right\}$
@@
When this is a singleton set $\{n\}$, we output $[n,n+1)$ and modify $M$ as before with $M\leftarrow\mbox{produce}(n,M)$; the bounds $I_{x}$ and $I_{y}$ do not change. Otherwise we leave $M$ unchanged and output upper and lower bounds on $M$, i.e. the min and max of ${\frac{axy+bx+cy+d}{exy+fx+gy+h}}$ subject to $x\in I_{x},y\in I_{y}$: denote this interval $\rho\left(\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right),I_{x},I_{y}\right)$. These bounds are easily computed, see section (3.5). To retain unlimited accuracy, the bounds must be computed with exact rational arithmetic.
@@
  {input $M$: bihomomorphic matrix}
  {output $z$: CF expansion of $M(x,y)$ represented as list of intervals}
@@
  $I_{x},I_{y}\leftarrow(-\infty,\infty),(-\infty,\infty)$current bounds on remaining parts of $x,y$}
@@
     if $x_{i}$ is ambiguous then
@@
        $M\leftarrow\mbox{ingest\_x}(x_{i},M)$
        $I_{x}\leftarrow[1,\infty)$
     end if
     if $y_{i}$ is ambiguous then
@@
     else
        $M\leftarrow\mbox{ingest\_y}(y_{i},M)$
        $I_{y}\leftarrow[1,\infty)$
@@
     if $\mathcal{R}(M,I_{x},I_{y})$ is not a singleton then
@@
     else
        $z_{j}\leftarrow\mathcal{R}(M,I_{x},I_{y})$
        $M\leftarrow\mbox{produce}(z_{i},M)$
@@
Figure 3. Algorithm for arithmetic on two CFs
3.4. Possible Optimizations of the CF Arithmetic Algorithm
3.4.1. Winnowing Interval Sequences
When we have to ingest the next interval from $x$ or $y$, and we see an ambiguous interval at the head of the list, it is not necessary to ingest that interval; we may skip it and go on to the next element of the list, which will give a tighter bound on the next term. Any such heuristic must guarantee that, given an infinite list of ambiguous input intervals, we eventually ingest something. Similarly, the algorithm does not need to output every ambiguous interval of $z$; but we must guarantee that we eventually output something, even if we are generating an infinite list of ambiguous intervals.
3.4.2. Taming Large Coefficients
If a high degree of accuracy is required, the coefficients of $M$ may grow very large; performing rational arithmetic on arbitrarily large integers will incur a performance penalty. We might be able to take advantage of the fact that we do not require the tightest possible bounds on the next term; the only requirement for correctness is that the bounds converge to the correct value. We might divide every coefficient of $M$ by some large integer $k$, rounding each result up or down in whichever direction slightly widens the range of $M$. We might do the same with intervals whose numerators and denominators are very large. To guarantee convergence, we must never make an adjustment which makes a new bound weaker than the current bound.
Any such decision to “simplify” $M$ must consider the gap between the upper and lower bounds of $M$; if the gap is large relative to the reciprocals of the coefficients, we might be carrying too many digits. Remember that, with lazy evaluation, all computations are driven by how much accuracy we need; if the gap is small, yet we are still continuing the computation, it means that we need even tighter bounds. In such cases large coefficients might be unavoidable.
3.5. Computing the range of $M$
To compute min and max of $M$ subject to $x_{\ell}<x<x_{u},y_{\ell}<y<y_{u}$:

### v2

@@ 3.1 Failure to Converge
$z_{0},z_{1},\cdots]={\frac{9xy+3x+3y+1}{4xy+2x+2y+1}}$ (4)
$z_{0},z_{1},\cdots]={\frac{49xy+21x+21y+9}{25xy+20x+10y+4}}$ (5)
$z_{0},z_{1},\cdots]={\frac{289xy+119x+119y+49}{144xy+60x+60y+25}}$ (6)
@@
The source of the problem is now clear: we should have $z=[z_{0},z_{1}]=[2,\infty]$, but no finite number of terms of $x$ and $y$ can ever suffice to determine that $z_{0}=2$. It is always possible that one input is in fact less than $\sqrt{2}$, so it is always possible that $z<2$ and $z_{0}=1$.
Our approach to the problem (see [6] for an entirely different approach) applies basic notions from interval arithmetic ([10, 8]). We separate the internal representation of a CF as a data structure from the mathematical notion of a sequence of integer terms. The internal representation will contain explicit terms (i.e. $z_{k}=n$) whenever such can be determined, but might also contain arbitrary bounds on the tail $z_{k},z_{k+1}\cdots]$, such as the sequence of bounds we derived for $\sqrt{2}*\sqrt{2}$. This is a direct generalization, because the explicit term $z_{k}=n$ is equivalent to the bound
@@
Therefore, when a CF is constructed explicitly from a known sequence of integer terms $a_{0},a_{1},\cdots]$, it will be represented internally as the sequence of half-open intervals
$a_{0},a_{0}+1),[a_{1},a_{1}+1),\cdots]\ .$
@@
At the opposite extreme, the internal representation for $(\sqrt{2})^{2}$ will consist entirely of ever-tighter bounds on the first term, i.e. the sequence of nested intervals
$\left[\left[\frac{16}{9},\frac{9}{4}\right),\left[\frac{49}{25},\frac{100}{49}\right),\left[\frac{576}{289},\frac{289}{144}\right),\cdots\right]\ .$
@@
The general idea (elaborated more fully in section 3.3) is that we will generate output on every iteration of the arithmetic algorithm; when we cannot produce the next term of $z$, we output upper and lower bounds on the current homographic matrix, leaving the matrix unchanged. Reading such output, there can be no ambiguity about the term to which a bound applies; the first bound applies to $z_{0},z_{1}\cdots]$, and as soon as we encounter an explicit term $[a,a+1)$ we know that $z_{k}=a$ and the next bound will apply to $z_{k+1},z_{k+2}\cdots]$. This implies that any finite prefix of the output sequence will determine a finite prefix $z_{0}\cdots z_{k-1}]$ of $z$, followed by a bound $\ell_{k},u_{k})$ on the tail $z_{k},z_{k+1}\cdots]$. The exact value of $z$ is between $z_{0}\cdots z_{k-1},\ell_{k}]$ and $z_{0}\cdots z_{k-1},u_{k}]$, if we extend the usual CF notation by allowing a rational number, instead of an integer, as the last term.
@@
Storage of a long sequence of intervals is not an issue for implementation; successive bounds on the same term are always tighter, so there is no need to retain earlier bounds when a new one is obtained.
A sequence of intervals will represent a real number if every ambiguous interval $I$ is immediately followed either by another ambiguous interval $I^{\prime}\subset I$, or by an integer term $a$ with $a\in I$; if the sequence ends with an infinite sequence of ambiguous intervals, the intersection of that infinite sequence must be a singleton $\{a\}$. We call such a sequence valid. It will be easy to verify that, if $x$ and $y$ are valid, the result of the arithmetic algorithm $z=M(x,y)$ is valid. The question of validity will become somewhat more subtle in section 4 when we consider transcendental functions of a CF.
3.2 Extracting Terms
@@
Note that our interval representation of CF arithmetic makes it impossible to simply ask a question like “what are the first five terms of z=xyz=xy?” An explicit list of terms can only be relative to a required degree of accuracy $\varepsilon$. Our arithmetic algorithms consume and return theoretically infinite sequences of intervals; extracting a finite sequence of terms is a separate “post-processing” step, carried out only when a numerical approximation is required. To obtain such an approximation, we find a prefix of the output interval sequence which is long enough to give
$\bigl{|}[z_{0},\cdots z_{k-1},\ell_{k}]-[z_{0},\cdots z_{k-1},u_{k}]\bigr{|}\leq\varepsilon\ .$
Then our rational approximation is $z_{0},\cdots z_{k-1},z^{\prime}_{k}]$, where $z^{\prime}_{k}$ is some integer in the interval $\ell_{k},u_{k})$. We might as well always take $z^{\prime}_{k}=\lfloor u_{k}\rfloor$, which is the only integer in the interval if the interval is small.
In the case of $z=(\sqrt{2})^{2}$, any $\varepsilon$ will yield the same result, $z=[2]$. More generally, whenever $M(\mathbf{x},\mathbf{y})=[z_{0},\cdots z_{k}]$ is rational, the rational approximation will eventually be $z_{0},\cdots z_{k}]$. As in section (2.1), this follows simply from the fact that upper and lower bounds on (3) will approach $z$ as bounds on $x$ and $y$ approach $\mathbf{x}$ and $\mathbf{y}$.
3.3 An Algorithm for CF Arithmetic
@@
We now present the full details of the algorithm suggested in the previous section. We augment the matrix $M$ with a pair of intervals $\{I_{x},I_{y}\}=\{[x_{\ell},x_{u}),[y_{\ell},y_{u})\}$, the current bounds on the remaining tails of the inputs $x,y$. These intervals start at $(-\infty,\infty)$, since the first term could be any integer.
When we read an interval of the form $[a,a+1)$ from $x$ or from $y$, we modify $M$ as before with $M\leftarrow\mbox{ingest\_x}(a,M)$ or $M\leftarrow\mbox{ingest\_y}(a,M)$, and we change that variable’s bound to $1,\infty)$. When we read an ambiguous interval $\ell,u)$ (i.e. a bound that does not determine the next term of the input), we replace the bound $I_{x}$ or $I_{y}$ with $\ell,u)$, leaving $M$ unchanged.
@@
When $M$ is unambiguous with $k=\lfloor M(x,y)\rfloor$, we output $[k,k+1)$ and modify $M$ as before with $M\leftarrow\mbox{produce}(k,M)$; the intervals $I_{x}$ and $I_{y}$ do not change. Otherwise we leave $M$ unchanged and output the interval from min to max of $M(x,y)$ subject to $x\in I_{x},y\in I_{y}$: denote this interval $\rho\left(M,I_{x},I_{y}\right)$. These bounds are easily computed, see section (3.4). To retain unlimited accuracy, the bounds must be computed with exact rational arithmetic.
3.3.1 Recursive Definition
We can express the algorithm as a recursive function, which is essentially the same as our Haskell implementation. We first define versions of $\mbox{ingest\_x},\mbox{ingest\_y}$ that consider the bounds on $x$ and $y$. Given an explicit term $[s,s+1)$,
$\mbox{ingBd\_x}([s,s+1),M,I_{x},I_{y})=(\mbox{ingest\_x}(s,M),(1,\infty),I_{y})$
while for an ambiguous interval $I_{0}$,
$\mbox{ingBd\_x}(I_{0},M,I_{x},I_{y})=(M,I_{0},I_{y})$
with a similar definition for ingBd_y.
We then define $\mbox{arith}(x,y,M,I_{x},I_{y})$ to be the representation of $M(x,y)$ as a list of intervals, where $x,y$ are lists of intervals and $M$ is bihomographic. We use the standard recursive notation $x=x_{0}\negthickspace:\negthickspace x_{s},y=y_{0}\negthickspace:\negthickspace y_{s}$ to denote the head and tail of a list888i.e. $a\negthickspace:\negthickspace[]=[a]$ and $a\negthickspace:\negthickspace[b,c,\cdots]=[a,b,c,\cdots]$. If $\rho(M,I_{x},I_{y})$ determines $a$ as the next output term, then
$\mbox{arith}(x,y,M,I_{x},I_{y})=(a,a+1)\negthickspace:\negthickspace\mbox{arith}(x,y,\mbox{prod}(a,M),I_{x},I_{y})$
while if $\rho(M,I_{x},I_{y})$ is ambiguous
$\mbox{arith}(x,y,M,I_{x},I_{y})=\rho(M,I_{x},I_{y})\negthickspace:\negthickspace\mbox{arith}(x_{s},y_{s},\mbox{ingBd\_y}(y_{0},(\mbox{ingBd\_x}(x_{0},M,I_{x},I_{y}))))$
If $M=\infty$ (i.e. the denominator is zero), we have exhausted a pair of rational inputs to produce a rational result, and can terminate the recursion with $\mbox{arith}(x,y,M,I_{x},I_{y})=[\infty]$.
@@ 3.3.1 Recursive Definition
 {input $M$: bihomographic matrix}
 {output $z$: CF expansion of $M(x,y)$, represented as a list of intervals}
@@
$I_{x},I_{y}\leftarrow(-\infty,\infty),(-\infty,\infty)$ {current bounds on remaining parts of x,yx,y}
@@
  if $x_{i}=[a,a+1)$ (where $a$ is an integer) then
   $M\leftarrow\mbox{ingest\_x}(a,M)$
   $I_{x}\leftarrow[1,\infty)$
  else
@@
  end if
  if $y_{i}=[a,a+1)$ (where $a$ is an integer) then
   $M\leftarrow\mbox{ingest\_y}(a,M)$
   $I_{y}\leftarrow[1,\infty)$
@@
  if $(M,I_{x},I_{y})$ is unambiguous with $\lfloor M(x,y)\rfloor=a$ for all $x\in I_{x},y\in I_{y}$ then
   $M\leftarrow\mbox{produce}(a,M)$
   $z_{j}\leftarrow[a,a+1)$
  else
@@
Figure 3: Algorithm for arithmetic on two CFs (without optimizations)
3.4 Computing the range of $M$
To compute $\rho\left(\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right),[x_{\ell},x_{u}),[y_{\ell},y_{u})\right)$: